OpenStudy (abymartinez15):

Can someone help me please A landscaper is selecting two trees to plant. He has five to choose from. Three of the five are deciduous and two are evergreen. What is the probability that he chooses trees of two different types? Express your answer as a percent.

OpenStudy (luhvon):

are there any numbers or percentages given for this problem?

OpenStudy (abymartinez15):

No

OpenStudy (luhvon):

im guessing u r suppose to use where they said the thing about the 3 and the 5 part as your percentage?

OpenStudy (abymartinez15):

Like i dont get it ? Wym

OpenStudy (welshfella):

Total number of combinations of 2 from 5 is 5C2 = (5*4)/(2*1) Hint: There are 2 ways he can pick one evergreen and 3 ways he can pick a deciduous.

OpenStudy (abymartinez15):

So after you multiply you divide 20/2 right?

OpenStudy (abymartinez15):

And it will be 10 ?

OpenStudy (welshfella):

yes 10 is all the possible combinations of 2 trees out of 5 and possible wyas to pick 2 different trees = 3*2 = 6

OpenStudy (abymartinez15):

So the answer would be 10 or 6?

OpenStudy (welshfella):

so required probability = number of possible ways to get different trees / total ways to pick 2 multiply this by 100 to convert to percentage.

OpenStudy (abymartinez15):

That would be 6 and 10?

OpenStudy (welshfella):

6 divided by 10

OpenStudy (abymartinez15):

60?

OpenStudy (welshfella):

60% yes

OpenStudy (welshfella):

6 -- * 100 10

OpenStudy (abymartinez15):

Thank you much

OpenStudy (welshfella):

yw